OpenStudy (anonymous):

What are the complex 5th roots of 5-5√3i?

OpenStudy (anonymous):

write in trig form first

OpenStudy (anonymous):

you know how to do that?

OpenStudy (anonymous):

Ummm. 10(cos60 - isin(60)

OpenStudy (anonymous):

not sure about the 10 l

OpenStudy (anonymous):

\[\sqrt{5^2+(5\sqrt3)^2}\]

OpenStudy (anonymous):

oh you win 10 but there is a mistake there

OpenStudy (anonymous):

it should be \[10\left(\cos(\theta)+i\sin(\theta)\right)\] not minus

OpenStudy (anonymous):

Gotcha

OpenStudy (anonymous):

angle is not 60

OpenStudy (anonymous):

dang draw tool is not working if you can draw it you will you are in quadrant 4, not 1

OpenStudy (anonymous):

Would it be... 300?

OpenStudy (anonymous):

yeah i guess if you are working in degrees

OpenStudy (anonymous):

now you want the fifth roots take \(\sqrt[5]{10}\) as the modulus, and divide the angle by 5 so the first answer will be \[\sqrt[5]{10}\left(\cos(60)+i\sin(60)\right)\]

OpenStudy (anonymous):

Okay...

OpenStudy (anonymous):

What about the other four?

OpenStudy (anonymous):

one way to do it is to divide the circle up in to five equal parts, with \(60\) degrees as one of them the other way is to keep adding 360 to the angle and divide by 5 again \[660\div 5=132\] so the next one is \[\sqrt[5]{10}\left(\cos(132)+i\sin(132)\right)\]

OpenStudy (anonymous):

lather, rinse, repeat

OpenStudy (anonymous):

Gotcha. Thanks!